OpenStudy (gahm8684):

What tha heck are they asking me? Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0 Numerical Answers Expected!

OpenStudy (gahm8684):

please help

OpenStudy (anonymous):

2(x+1)(2x+1)

OpenStudy (gahm8684):

what is that

OpenStudy (gahm8684):

thats equals to 4x2 + 6x + 2

OpenStudy (anonymous):

oh i see hold on

OpenStudy (anonymous):

Are they asking you to find the the value for -b, b^2-4ac, and 2a?

OpenStudy (gahm8684):

no idea what they are asking me i got the same exact question

OpenStudy (gahm8684):

lol

OpenStudy (anonymous):

Well if they're asking for the values, it's -b=-6, b2-4ac=4, and 2a=8

OpenStudy (gahm8684):

so do i just put 8

OpenStudy (gahm8684):

you were right

OpenStudy (gahm8684):

i just had to put a number such as 4

OpenStudy (gahm8684):

i put b2-4ac=4 in other words i put just 4